3.19 Kelle C offices. Peter Kelle h time it took to complete several recent jobs. So are very unreliable. A list of activities and their op pletion time, the completion time (all in days) for a new contract are given in following table: arpet and Trim installs carpet in commercial as been very concerned with the amount of com- most likely completion time, and the pessimistic timistic the
media%2Fa5f%2Fa5ffb7bb-eee1-413f-9f1f-bb
Answer #1

a)

Activi Optimistic(a Most likely(m) Pessimistic(b Expected time (B2+4*C2+D2)/6 B3+4 C3+D3)/6 (B4+4*C4+D4/6 (B5+4*C5+D5)/6 (B6+1 Activity Optimistic(a) Most likelv(m) Pessimistic(b) Expected time Variance 0.6944 5.83 3.67 2.00 7.00 4.00 10.00 2.17 6.00

b)

13 oat 36-33 9 3623 7) End 6.33 368.3

The total project completion time is 36.33

Critical path is C-D-E-F-H-K

c)

3 0808 18 | 0 2 3 929-13 2 75029 4 37 733 1-33 29-331 686 213 3 1 112 A B C D E F G H 456789d)

First determine Z

Z= (X-Tcp)/Stdcp

Tcp- Time along critical path

Stdcp- Standard deviation along critical path

Vcp- Variance along critical path= Sum of variances of critical activities

Vcp = 5.2222

Stdcp = squareroot of Vcp = sqrt(5.2222) = 2.2852

Z = (40-36.33)/2.2852 = 1.606

Probability is 0.9459